# SCaLAR NITK at SemEval-2024 Task 5: Towards Unsupervised Question Answering system with Multi-level Summarization for Legal Text

### v1

@@ 1 Introduction
The Domain of Law demands sheer expertise and experience for a human to master, but it takes much more to teach a machine the same. Legal NLP (Zhong et al., 2020) is advancing at a rapid pace, and the advent of Transformers Vaswani et al. (2017) has widened the prospects of research in this area. However, the intricate nature of Legal Texts and the underlying complex relationships between entities make it difficult even for state-of-the-art Language models like BERT Devlin et al. (2019) to capture the details effectively. To advance our understanding of the reasoning ability of LLMs in the legal domain, task 5 of SemEval-2024 was proposed Bongard et al. (2022). The objective of this task is to discern the accurate responses to legal inquiries in U.S. Civil Procedure, as posited by the organizers. The questions and answers adhere to a Multiple-choice question-answering model, with accompanying explanations provided to facilitate comprehension of the legal concepts associated with each question. We have also released the code on GitHub  [^1] https://github.com/haricharan189/Semevaltask5.
@@ 2 Background
Field
Text
Explanation
The most basic point to understand about supplemental jurisdiction …….. on this basic purpose of Ârticle 1367(a).
Question
This and that. Garabedian, ……….. are treated fairly.
Answer
has constitutional authority …………… under Ârticle 1367(a).
Label
0
Analysis
Here, the Ârticle 1983 claim ………… Amendment claim.
Complete analysis
This is pretty straightforward ………… D is the best choice here.
@@ 4.2.1 Word2Vec-Cosine system
[…]
@@ 5.2 Unsupervised Models
Word2Vec and GloVe embeddings were both generated with an embedding size of 5. However, there were differences in the window length used during training: for Word2Vec embeddings, a window length of 7 was utilized, while GloVe embeddings were trained with a window length of 10. In the case of GloVe, the training process spanned 30 epochs, employing a learning rate of 0.05 to optimize the model parameters.
@@ 6 Results
Baseline
0.798 0.444 0.7449 0.4269
Multi-level approach
0.74 0.65 0.4898 0.4102
Multi-Feature approach
0.81 0.66 0.6224 0.4966
Word2vec-cosine
0.71 0.62 0.6429 0.5238
@@
GloVE-cosine 0.64 0.52 0.6020 0.4694
Transformer-cosine
0.60 0.46 0.5612 0.4150
@@ 7 Conclusion and Future scope
Future endeavors entail amalgamating these models into a unified super model. This super model would aggregate predictions from various models to yield a singular final prediction, enhancing overall performance and addressing the limitations of individual approaches. An alternative strategy involves leveraging Siamese networks to learn similarity, addressing challenges encountered by unsupervised models when all answers for a particular question are labeled as incorrect (0). By employing Siamese networks, we believe that the model can effectively capture nuanced similarities between question-answer pairs, and provide better predictions.

### v2

@@ 1 Introduction
The Domain of Law demands sheer expertise and experience for a human to master, but it takes much more to teach a machine the same. Legal NLP (Zhong et al., 2020) is advancing at a rapid pace, and the advent of Transformers Vaswani et al. (2017) has widened the prospects of research in this area. However, the intricate nature of Legal Texts and the underlying complex relationships between entities make it difficult even for state-of-the-art Language models like BERT Devlin et al. (2019) to capture the details effectively. To advance our understanding of the reasoning ability of LLMs in the legal domain Bongard et al. (2022), task 5 of SemEval-2024 was proposed Held and Habernal (2024). The objective of this task is to discern the accurate responses to legal inquiries in U.S. Civil Procedure, as posited by the organizers. The questions and answers adhere to a Multiple-choice question-answering model, with accompanying explanations provided to facilitate comprehension of the legal concepts associated with each question. We have also released the code on GitHub  [^1] https://github.com/haricharan189/Semevaltask5.
@@ 2 Background
Field Text
Explanation The most basic point to understand about supplemental jurisdiction …….. on this basic purpose of Ârticle 1367(a).
Question This and that. Garabedian, ……….. are treated fairly.
Answer has constitutional authority …………… under Ârticle 1367(a).
Label 0
Analysis Here, the Ârticle 1983 claim ………… Amendment claim.
Complete analysis This is pretty straightforward ………… D is the best choice here.
@@ 4.2.1 Word2Vec-Cosine system
      if $\lvert\text{similarity}[max\_id]-\text{similarity}[second\_max\_id]\rvert\leq 0.0005$ then
@@ 5.2 Unsupervised Models
Word2Vec and GloVe embeddings were both generated with an embedding size of 5. However, there were differences in the window length used during training: for Word2Vec embeddings, a window length of 7 was utilized, while GloVe embeddings were trained with a window length of 10. In the case of GloVe, the training process spanned 30 epochs, employing a learning rate of 0.05 to optimize the model parameters. These values of hyper-parameters were arrived after experimentation with several other values.
@@ 6 Results
Baseline 0.798 0.444 0.7449 0.4269
Multi-level approach 0.74 0.65 0.4898 0.4102
Multi-Feature approach 0.81 0.66 0.6224 0.4966
Word2vec-cosine 0.71 0.62 0.6429 0.5238
@@
GloVE-cosine 0.64 0.56 0.6020 0.4694
Transformer-cosine 0.60 0.46 0.5612 0.4150
@@ 7 Conclusion and Future scope
Future endeavors entail amalgamating these models into a unified super model. This super model would aggregate predictions from various models to yield a singular final prediction, enhancing overall performance and addressing the limitations of individual approaches. An alternative strategy involves leveraging Siamese networks to learn similarity, addressing challenges encountered by unsupervised models when all answers for a particular question are labeled as incorrect (0). By employing Siamese networks, we believe that the model can effectively capture nuanced similarities between question-answer pairs, and provide better predictions. Exploring other kind of summarizers and using other transformers for summarization such BART Lewis et al. (2020) may also increase the overall performance of all the systems used in this paper. Data augmentation Feng et al. (2021) can also be implemented to get better Word2Vec and GloVE embeddings.
Acknowledgements
We would like to thank the organizers, reviewers and SemEval - 2024 Chairs for their valuable insights and helpful suggestions.